# Codebook-enabled Generative End-to-end Semantic Communication Powered by Transformer

### 22 Jan 2024 v1

@@ I Introduction
In recent years, the development of Transformer [10] has been widely recognized, particularly with the emergence of large models such as ChatGPT, which has propelled artificial intelligence to new heights. The self-attention mechanism employed by Transformer allows each position in the input sequence to dynamically attend to all other positions, enabling the model to directly gather information from the entire input sequence. This emphasis on global information over local details endows Transformer with outstanding performance in various tasks, especially in situations that require consideration of long-range dependencies. In [11], Zhou et al. addressed the matching problem between degraded image feature vector and high-quality image feature vector by introducing the Transformer. Inspired by their work, this paper incorporates the Transformer to tackle the challenge of recovering image feature map, as transmitted information, after being contaminated by noise. The goal is to restore these features to the state of indices suitable for image reconstruction.
@@ III-A Stage 1: Joint Construction of Semantic Codec and Codebook
Training Objectives. In [11] [12], for the E2E network in Stage 1, based on the network’s input and output, we can establish a result-based loss function $\mathcal{L}_{result}$:
@@ V Conclusion
This paper proposes a robust codebook-assisted image semantic communication method, via transmitting the feature map directly to the receiver, and leveraging Transformer to predict and correct feature map with errors based on the global information of the feature maps and the assistance of codebook, achieving excellent source image reconstruction. The use of a codebook significantly enhances the performance of vector-to-index Transformer, thereby improving communication stability and efficiency. Furthermore, images generated based on high-quality map from the learned codebook surpass those of comparative methods in terms of visual aesthetics. Additionally, we conduct a thorough analysis of the characteristics of generative semantic communication in image transmission, elucidating the superiority of generative semantic communication methods. This work further substantiates the promising prospects and value of generative network structures in the field of semantic communication.

### 05 Mar 2024 v2

@@ I Introduction
In recent years, the development of Transformer [10] has been widely recognized, particularly with the emergence of large models such as ChatGPT, which has propelled artificial intelligence to new heights. The self-attention mechanism employed by Transformer allows each position in the input sequence to dynamically attend to all other positions, enabling the model to directly gather information from the entire input sequence. This emphasis on global information over local details endows Transformer with outstanding performance in various tasks, especially in situations that require consideration of long-range dependencies. In [11], Zhou et al. addressed the matching problem between degraded image feature vector and high-quality image feature vector by introducing the Transformer. Inspired by their work [11] [12], this paper incorporates the Transformer to tackle the challenge of recovering image feature map, as transmitted information, after being contaminated by noise. The goal is to restore these features to the state of indices suitable for image reconstruction.
@@ III-A Stage 1: Joint Construction of Semantic Codec and Codebook
Training Objectives. For the E2E network in Stage 1, based on the network’s input and output, we can establish a result-based loss function $\mathcal{L}_{result}$:
@@ V Conclusion
This paper proposes a robust codebook-assisted image semantic communication method, via transmitting the feature map directly to the receiver, and leveraging Transformer to predict and correct feature map with errors based on the global information of the feature map and the assistance of codebook, achieving excellent source image reconstruction. The use of a codebook significantly enhances the performance of vector-to-index Transformer, thereby improving communication stability and efficiency. Furthermore, images generated based on high-quality map from the learned codebook surpass those of comparative methods in terms of visual aesthetics. Additionally, we conduct a thorough analysis of the characteristics of generative semantic communication in image transmission, elucidating the superiority of generative semantic communication methods. This work further substantiates the promising prospects and value of generative network structures in the field of semantic communication.